OpenStudy (anonymous):

A rectangular field is 40 yards long. If you ran diagonally across the field, you would run 50 yards. How many yards wide is the field?

OpenStudy (anonymous):

 Try solving for x using the Phythagorean Theorem, as we have a right triangle here.

OpenStudy (anonymous):

a^2+b^2=c^2 so... 40^2 + 50^2= c^2 therefore... 1,600+2,500= 4,100^2 take the square root of 4,100 and your answer would be.... .....64.03

OpenStudy (anonymous):